OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). (2 points) (A) y = 2x + 6 (B) y = − 1/2 x + 3 (C) y = − 1/2 x + 6 (D) y = 2x + 3 Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). (2 points) (A) y = 2x + 6 (B) y = − 1/2 x + 3 (C) y = − 1/2 x + 6 (D) y = 2x + 3 @Mathematics

OpenStudy (aroub):

y=mx+b , m=-1/2 , (6,3) 3=-1/2(6)+b 3=-3+b 6=b y=-1/2x+6 so it's (c)

OpenStudy (anonymous):

i thought soo thanks

OpenStudy (aroub):

Ok, no problem :)